# H2EAL: Hybrid-Bonding Architecture with Hybrid Sparse Attention for Efficient Long-Context LLM Inference

### v1

@@ V-C1 LongBench Result
Table IV demonstrates that H2EAL achieves comparable or even improved performance compared to full attention across the majority of datasets. The reason why sparse attention outperforms full attention is likely due to the ability to mitigate attention noise for long contexts. Figure 12 demonstrates the comparison between H2EAL and recent works [22, 16], and H2EAL consistently achieves higher accuracy than StreamingLLM [22] and H2O [16] across different datasets.
@@ ACKNOWLEDGMENTS
This work was supported in part by NSFC under Grant 62495102, Grant 92464104, Grant 62125401, and Grant U24A20287, in part by Beijing Outstanding Young Scientist Program under Grant JWZQ20240101004, in part by Beijing Municipal Science and Technology Program under Grant Z241100004224015, and in part by 111 Project under Grant B18001. We thank the authors of DuoAttention[19], Lserve[27] for releasing their code and for helpful discussions, which facilitated our implementation.

### v2

@@ V-C1 LongBench Result
Table IV demonstrates that H2EAL achieves comparable or even improved performance compared to full attention across the majority of datasets. The reason why sparse attention outperforms full attention is likely due to the ability to mitigate attention noise for long contexts. Figure 12 demonstrates the comparison between H2EAL and recent works [Xiao2023EfficientSL, Zhang2023H2OHO], and H2EAL consistently achieves higher accuracy than StreamingLLM [Xiao2023EfficientSL] and H2O [Zhang2023H2OHO] across different datasets.
@@ ACKNOWLEDGMENTS
This work was supported in part by NSFC under Grant 62495102, Grant 92464104, Grant 62125401, and Grant U24A20287, in part by Beijing Outstanding Young Scientist Program under Grant JWZQ20240101004, in part by Beijing Municipal Science and Technology Program under Grant Z241100004224015, and in part by 111 Project under Grant B18001. We thank the authors of DuoAttention[Xiao2024DuoAttentionEL], Lserve[Yang2025LServeEL] for releasing their code and for helpful discussions, which facilitated our implementation.